OpenStudy (anonymous):

cos^2(5pi/12). I know that is (1+ cos5pi /12) 2 , but am stuck after that

OpenStudy (lgbasallote):

actually... \[\Large \cos^2 (\frac{5\pi}{12}) \implies \left[\cos (\frac{5\pi}{12})\right]^2\] i dont think those are the same...

OpenStudy (anonymous):

i have to use the half angle formula

OpenStudy (anonymous):

right, as \(\frac{5}{12}\) is half of \(\frac{5}{6}\) so use the half angle formula with \(\theta=\frac{5\pi}{6}\)

OpenStudy (anonymous):

your mistake is in your post is that you need \(\theta =\frac{5\pi}{6}\) in the formula, not \(\frac{5\pi}{12}\)

OpenStudy (anonymous):

i.e. it should be \[\frac{1+\cos(\frac{5\pi}{6})}{2}\]

OpenStudy (anonymous):

oh i see